2) The present value of an annuity of $8,100 each year for eighteen years, assuming an opportunity cost of 11 percent is
Answer #1

Present value of Annuity is given by : -

\large Present Value of Annuity = Cash flow * [\frac{1-\frac{1}{(1+r)^{n}}}{r}]

  \large = 8100 * [\frac{1-\frac{1}{(1+0.11)^{18}}}{0.11}]

= \large 8100 *\frac{0.84718}{0.11}

= $62,383

Preset Value = $ 62,383

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.